Two cars leave the same place at the same time. The first drives in a straight line N 35 Degrees W at 30 miles per hour and the second drives in a straight line N 12 Degrees E at 40 miles per hour. After 1 hour, how far apart are the cars(to the nearest mile)?
1
Expert's answer
2010-12-05T05:55:09-0500
The distance between the cars equals to the side of triangle with other two sides of 30 and 40 miles and the


angle between them of (35+12)= 47 degrees.
According the Law of cosines:
C2 = a2+b2 - 2 ab cos(theta)
C2 = 302+402-2*30*40*cos(47) = 863.2
C = 29.4 miles.